### v1

# compat=1.1.0

Abstract
We develop a unified framework connecting quantum Tsallis statistics to electronic transport in strongly interacting systems. Starting from Rényi and Tsallis entropies, we construct a quantum Tsallis distribution that reduces to the conventional Fermi–Dirac distribution when $q=1$. For $q$ slightly deviating from unity, the correction term in the occupation function can be mapped to a $q$-deformed Schwarzian action, corresponding to soft reparametrization modes. Coupling these soft modes to electrons via the Fermi Golden Rule yields a modified scattering rate, which reproduces conventional Fermi-liquid behavior at low temperatures and linear-in-temperature resistivity at high temperatures. Using the memory matrix formalism, we analyze magnetotransport, finding a linear-in-field magnetoresistance and a Hall angle consistent with Anderson’s two-lifetime scenario. At sufficiently low temperatures, both magnetoresistance and Hall response smoothly recover Fermi-liquid quadratic behaviors. This approach provides a controlled interpolation between Fermi-liquid and non-Fermi-liquid regimes, quantitatively linking $q$-deformation, soft-mode dynamics, and experimentally measurable transport coefficients in strange metals.
Introduction– Landau’s Fermi liquid theory has long served as the foundational framework for understanding conventional metals, wherein low-energy excitations are described by long-lived quasiparticles with renormalized effective masses. In this picture, electrical resistivity saturates at both low and high temperatures—due to phonon freezing and the Mott–Ioffe–Regel (MIR) bound, respectively [1, 2]. However, in a growing class of materials known as strange metals, the resistivity remains linear in temperature over a remarkably wide range without saturation, even beyond the MIR limit, indicating the breakdown of the quasiparticle description and the emergence of collective, nonlocal charge-carrying degrees of freedom [3, 4, 5, 6, 7, 8]. This non-Fermi liquid behavior in strange metals is typically characterized by linear-in-temperature ($T$-linear) resistivity and linear-in-magnetic-field ($B$-linear) magnetoresistivity, alongside scale-invariant dissipation governed by the Planckian time $\tau_{P}=\hbar/(k_{B}T)$ [9, 11, 12, 10, 13, 14]. Interestingly, the Hall angle in the cuprates exhibits a distinct scaling $\theta_{H}\sim 1/T^{2}$, resembling Fermi-liquid-like behavior and therefore standing in sharp contrast to the ubiquitous $T$-linear resistivity of strange metals [9, 10]. Experimental studies — notably the early work by Chien et al. — observed such Hall-angle scaling and its sensitivity to impurity scattering, underscoring the inadequacy of single-relaxation-time transport models and pointing instead to multiple, possibly channel-dependent, scattering mechanisms underlying strange-metal phenomenology [15].
In gravitational physics, a similarly radical departure from standard local descriptions arises in the resolution of the black hole information loss paradox via the island prescription [16, 17, 18, 19, 20, 21, 22]. In the Euclidean gravitational path integral, two dominant saddles contribute to the entanglement entropy of Hawking radiation: the “Hawking saddle”—which governs early times and leads to a monotonically increasing entropy—and the “replica wormhole saddle”—which dominates at late times and produces entropy saturation consistent with unitarity. The von Neumann entanglement entropy is computed using the replica trick,
$S=\lim_{q\to 1}S_{q}^{\mathrm{R\acute{e}nyi}}=\lim_{q\to 1}\frac{1}{1-q}\log\mathrm{Tr}\,\rho^{q},$ (1)
where $\rho$ is the reduced density matrix of the Hawking radiation. Geometrically, replica wormholes connect multiple copies of the spacetime, linking the island region with the radiation region in such a way that their mutual entanglement is effectively canceled. This nonlocal connection underlies the unitarity of the full evaporation process.
The retrieval of quantum information in gravity thus depends crucially on nonlocal saddle-point contributions. An intriguing question is whether analogous ideas can be applied to strongly correlated condensed matter systems such as strange metals, where transport deviates strongly from Fermi liquid theory and scale-invariant features suggest emergent nonlocal dynamics [23, 24]. We propose that the Tsallis entropy [25, 26], a nonextensive generalization of Boltzmann–Gibbs entropy, provides a natural framework for such an extension. The Tsallis parameter $q$ encodes deviations from extensivity, modeling systems with long-range correlations and power-law distributions.
From this viewpoint, the standard Fermi–Dirac distribution—and the corresponding Fermi liquid behavior—is analogous to the Hawking saddle: a disconnected, local description valid in regimes where quasiparticles dominate. Allowing the replica index $q$ to deviate slightly from unity, $q=1+\delta(T)$ with $\delta(T)=T/T_{0}\ll 1$ temperature dependent, yields a Tsallis-deformed fermionic distribution that encodes non-Fermi liquid behavior, reminiscent of the contribution of replica wormholes in semiclassical gravity. Such a deformation can be captured by a Schwarzian effective action [27, 28], which governs the soft mode dynamics of nearly-$\mathrm{AdS}_{2}$ geometries. This correspondence suggests a dual statistical picture: the strange metal phase, with its scale-invariant transport and Planckian dissipation, may be understood as the condensed matter analogue of gravitational systems dominated by wormhole saddles. In the following, we develop this idea by deriving the quantum Tsallis distribution, relating it to both Rényi and von Neumann entropies, and showing how its low-energy limit maps to the Schwarzian theory that appears in the strongly interacting quantum matter.
Quantum Tsallis statistics- For a quantum system described by a density matrix $\rho$, the Tsallis entropy of order $q\in\mathbb{R}$ is defined as
[…]
The Tsallis entropy is closely related to the Rényi entropy, defined in equation (1) and the two obey the following relation
[…]
This expression shows that the Tsallis entropy provides a power-law expansion of the exponential Rényi entropy. In the limit $q\to 1$, both Tsallis and Rényi entropies smoothly reduce to the von Neumann entropy,
[…]
The Tsallis entropy is non-additive for independent systems, reflecting possible long-range correlations
$S_{q}^{\mathrm{Tsallis}}(\rho_{A}\otimes\rho_{B})=S_{q}^{\mathrm{Tsallis}}(\rho_{A})+S_{q}^{\mathrm{Tsallis}}(\rho_{B})+(1-q)\,S_{q}^{\mathrm{Tsallis}}(\rho_{A})\,S_{q}^{\mathrm{Tsallis}}(\rho_{B}).$
From the Tsallis and Rényi entropies, we can obtain the corresponding distribution. Using the Langrangian multiplier method with constraints $Tr\rho=1$ and ${\rm Tr}(\rho H)=\langle E\rangle$, we can write down the Langrangian as follows
$\mathcal{L}=\frac{1}{1-q}\ln\bigg({\rm Tr}\rho^{q}\bigg{missing})-\alpha({\rm Tr}\rho-1)-\beta\bigg{(}{\rm Tr}(\rho H)-\langle E\rangle\bigg{)}.$ (4)
We then obtain
$\delta\mathcal{L}=\frac{q}{1-q}\frac{{\rm Tr}(\rho^{q-1}\delta\rho)}{{\rm Tr}(\rho^{q})}-\alpha{\rm Tr}\delta\rho-\beta{\rm Tr}(H\delta\rho).$ (5)
From maximal entropy $\delta\mathcal{L}=0$ and after some manipulation, we find
$\rho\;=\;\frac{1}{Z_{q}}\,\Big{[}\,1-(1-q)\beta H\,\Big{]}_{+}^{\frac{1}{1-q}}\;,$
[…]
which is finite whenever the trace exists; for unbounded $H$, one either has a high-energy cutoff for $q<1$ (since $1-(1-q)\beta\varepsilon\geq 0$ implies $\varepsilon\leq 1/[(1-q)\beta]$) or a power-law tail for $q>1$ with convergence controlled by the growth of the spectral density. In the Gibbs limit $q\to 1$, the well-known identity $\lim_{q\to 1}\big{[}1-(1-q)\beta H\big{]}^{1/(q-1)}=e^{-\beta H}$ applies eigenvalue-wise and gives $\rho\;\to\;\frac{e^{-\beta H}}{{\rm Tr}(e^{-\beta H})},\qquad Z_{q}\;\to\;Z={\rm Tr}(e^{-\beta H}),$ so the Tsallis ensemble continuously deforms the canonical ensemble and recovers it at $q=1$.
For quantum statistics with indistinguishable particles, the same $q$-deformation propagates to the grand-canonical occupation numbers once one replaces the ordinary exponential by the $q$-exponential while enforcing the exchange symmetry. The generalized quantum occupations can be written compactly as
$n_{q}(\varepsilon)\;=\;\frac{1}{\big{[}\,1-(1-q)\beta(\varepsilon-\mu)\,\big{]}_{+}^{\frac{1}{1-q}}+a},$
where $\beta=1/T$ and $\mu$ is the chemical potential; the parameter $a=+1$ gives the Tsallis–Fermi–Dirac case (Pauli exclusion), a=1a=-1 the Tsallis–Bose–Einstein case (Bose enhancement), and $a=0$ the Tsallis–Boltzmann classical limit. This expression reduces to the standard Fermi–Dirac/Bose–Einstein/Boltzmann forms as $q\to 1$. If one prefers the alternative but equivalent representation used in some conventions, the denominator can be written as $\big{[}1+(q-1)\beta(\varepsilon-\mu)\big{]}^{1/(q-1)}+a$, noting that $1+(1-q)x]^{1/(q-1)}=\big{(}[1+(q-1)(-x)]^{1/(q-1)}\big{)}^{-1}$; either choice is consistent with the operator formula for $\rho$ above, provided the same convention for the $q$-exponential is used throughout and the spectral positive-part enforces positivity and, when $q<1$, the finite-energy support of the ensemble.
q-deformed Schwarzian action- We show that the Tsallis-modified fermion distribution can be interpreted as arising from an effective reparametrization theory for the low-energy modes of nearly-$\mathrm{AdS}_{2}$ gravity. In the Tsallis framework, deviations from the Fermi-Dirac (FD) distribution are parameterized by (q1)(q-1), which controls the strength of non-extensive effects. At $q=1$, the standard FD form is recovered. For $q\neq 1$, the distribution acquires a correction term $\phi(x)$ that is quadratic in the dimensionless energy variable $x$ and weighted by the thermal factor $n_{\mathrm{FD}}(x)[1-n_{\mathrm{FD}}(x)]$. We reinterpret $x$ as the Euclidean frequency operator, $x\leftrightarrow-i\partial_{\tau}$, and promote $\phi(x)$ to a local functional $\Phi[f](\tau)$ of the time-reparametrization mode $f(\tau)$. This functional encodes the same fluctuation structure as $\phi(x)$, but now in the geometric language of boundary diffeomorphisms. In the $q\to 1$ limit, $\phi(x)$ vanishes and $\Phi[f](\tau)$ disappears, leaving only the Schwarzian sector; this ensures a smooth interpolation between the Tsallis-deformed and undeformed cases.
We start from the expansion of the quantum Tsallis distribution around $q=1$:
$n_{q}(x)=n_{\mathrm{FD}}(x)+(q-1)\,\phi(x)+\cdots,\qquad x=\beta(\epsilon-\mu),$ (6)
where
$\phi(x)=\frac{x^{2}}{2}\,n_{\mathrm{FD}}(x)\big{[}1-n_{\mathrm{FD}}(x)\big{]}.$ (7)
Thus, $\phi(x)$ can be interpreted as the variance of the Fermi–Dirac occupation number, modulated by a quadratic prefactor $x^{2}$. Since $n_{\mathrm{FD}}(1-n_{\mathrm{FD}})$ is maximized at $x=0$ and decays away from the Fermi surface, this thermal weight confines the correction to states near the Fermi level. In the frequency–time correspondence (taking the Euclidean Matsubara convention), we have $x=\beta\omega$ and $i\partial_{\tau}\leftrightarrow\omega$. Hence,
$x^{2}\ \longleftrightarrow\ \beta^{2}\omega^{2}\ \longleftrightarrow\ -\beta^{2}\partial_{\tau}^{2}.$ (8)
This indicates that the quadratic factor $x^{2}$ corresponds to the $\omega^{2}$ contribution in frequency space, which in the time domain is equivalent to a second-order Euclidean time derivative (with an overall sign depending on Fourier conventions). Under the assumption of low-frequency dynamics and a sufficiently smooth response kernel, the frequency-domain correction $\phi(x)$ can be promoted to a local time-domain functional $\Phi[f](\tau)$. In this sense, the (q1)(q-1) correction to the distribution can be reinterpreted as a soft-mode contribution involving second-order time derivatives of the reparametrization mode $f(\tau)$. This identification is approximate: for strongly nonlocal frequency responses $|x|\gg 1$, the local replacement by $\partial_{\tau}^{2}$ is no longer valid. Motivated by the structure of $\phi(x)$, we define the local functional
$\Phi[f](\tau)=\big{(}u^{\prime}(\tau)\big{)}^{2}-2\,u^{\prime\prime}(\tau),\qquad u(\tau)=\log f^{\prime}(\tau),$
where $\left(\frac{f^{\prime\prime}}{f^{\prime}}\right)^{2}=(u^{\prime})^{2}$ and $\partial_{\tau}\log f^{\prime}(\tau)=u^{\prime}(\tau)$. The quadratic term in $\phi(x)$ corresponds to $(u^{\prime})^{2}$ in the geometric language, while the remaining structure naturally maps to the second derivative term $u^{\prime\prime}(\tau)$. The $q$-deformed effective action is then schematically written as
$S_{q}[f]=-C\int\{f,\tau\}\,+\,(q-1)\,\lambda\int\Phi[f](\tau)\,+\,\mathcal{O}((q-1)^{2}),$
where $\{f,\tau\}$ is the Schwarzian derivative and $\lambda$ is a coupling determined by the microscopic model. For small deviations from the identity reparametrization, $f(\tau)=\tau+\varepsilon(\tau)$, the quadratic effective action for the soft-mode fluctuations takes the form
$S_{q}^{(2)}[\varepsilon]=\frac{\bar{C}_{\rm eff}}{2}\int_{0}^{\beta}d\tau\,(\varepsilon^{\prime\prime}(\tau))^{2},$ (9)
where the effective coupling is renormalized by the $q$-deformation,
$\bar{C}_{\rm eff}=\bar{C}+\delta C\,(q-1)+\mathcal{O}((q-1)^{2}).$
In the limit $q\to 1$, one recovers the standard Schwarzian coupling $\bar{C}$ and the familiar zero-mode solutions
$f(\tau)=\frac{a\tau+b}{c\tau+d},\qquad ad-bc\neq 0,$
up to small fluctuations of the soft mode $\varepsilon(\tau)$. The quadratic action and its solution thus provide the leading approximation for small q1q-1 deformations. For $q\to 1$, the deformation term vanishes and the partition function describes the motion of a particle along the boundary of $\mathrm{AdS}_{2}$ with configuration space $SL(2,\mathbb{R})/U(1)$. Incorporating the Tsallis correction, the $q$-deformed Euclidean partition function takes the form
$Z_{q}=\int\frac{\mathcal{D}f}{SL(2,\mathbb{R})}\exp\left\{-\bar{C}\int_{0}^{\beta}d\tau\left[\{f(\tau),\tau\}-(q-1)\left(\frac{f^{\prime\prime}(\tau)}{f^{\prime}(\tau)}\right)^{2}\right]\right\},$ (10)
where $\bar{C}$ is the Schwarzian coupling constant. The $q$-deformation of Eq. (10) contributes an additional quadratic term to the effective action, which can be interpreted as a renormalization of the Schwarzian stiffness and modifies the dynamics of the soft reparametrization mode. The corresponding effective Hamiltonian reads
$H=-\frac{1}{2\bar{C}}\frac{d^{2}}{du^{2}}+\frac{1}{8\bar{C}\sinh^{2}(u)},$ (11)
where $u$ is the geodesic distance on the group manifold, and $u\to\infty$ corresponds to asymptotically free motion. The Hilbert space is given by the continuous unitary series of $SL(2,\mathbb{R})$. The corresponding Schrödinger equation reads
$\left[-\frac{d^{2}}{du^{2}}+\frac{1}{4\sinh^{2}(u)}\right]\phi(u)=E\,\phi(u),$ (12)
with eigenfunctions
$\phi_{E}(u)=\sqrt{\sinh u}\,P^{-1/2+ik}_{1/2}(\cosh u),\qquad E=k^{2}.$ (13)
The spectral density is determined by the Plancherel measure of $SL(2,\mathbb{R})$ representations  [27, 28, 31, 32, 33]
$\rho(E)=\frac{\bar{C}}{2\pi^{2}}\sinh\!\left(2\pi\sqrt{2\bar{C}E}\right),$ (14)
with asymptotic behaviors
$\rho(E)\approx\frac{\bar{C}^{3/2}}{\pi}\sqrt{2E},~E\ll 1;~~~\rho(E)\sim\frac{\bar{C}}{4\pi^{2}}\exp\!\left(2\pi\sqrt{2\bar{C}E}\right),~E\gg 1.$ (15)
Equation (15) exhibits the familiar square-root scaling of the density of states at low energies, analogous to the $1D$ density of states in a quantum well. In the high-energy limit $E\gg 1$, the spectral density grows exponentially, which is consistent with a Cardy-like behavior [34]. This exponential growth reflects the rapid proliferation of high-energy states in the system, which can be interpreted as a manifestation of a large number of chaotic degrees of freedom [35, 36]. In the context of the Sachdev-Ye-Kitaev (SYK) model and its dual nearly-$AdS_{2}$ gravity description, this corresponds to the presence of strongly interacting modes whose dynamics are effectively governed by the Schwarzian action. The Plancherel measure thus provides a precise connection between the group-theoretical structure of $SL(2,\mathbb{R})$ and the statistical distribution of energy levels in these quantum chaotic systems [37, 38].
Electron–Schwarzian Soft-Mode Coupling- We expand reparametrizations around the thermal saddle $f(\tau)=\tau+\varepsilon(\tau)$ and fix the $SL(2,\mathbb{R})$ zero modes. Writing Fourier components
$\varepsilon(\tau)=\sum_{n\in\mathbb{Z}}\varepsilon_{n}\,e^{-i\omega_{n}\tau},\qquad\omega_{n}=2\pi nT,\;n\neq 0,\pm 1,$
the quadratic Schwarzian action takes the form
$S^{(2)}_{\rm Sch}=2\pi\bar{C}T\sum_{n\geq 2}n(n^{2}-1)\,|\varepsilon_{n}|^{2},\qquad\varepsilon_{-n}=\varepsilon_{n}^{*}.$ (16)
The modes $n=0,\pm 1$ are excluded because they generate the $SL(2,\mathbb{R})$ transformations (constant shift, scaling, and inversion) under which the Schwarzian action is invariant. The $q$-deformation contributes an additional quadratic correction,
$\Delta S^{(2)}_{q}=2\pi\bar{C}T\,\gamma_{q}\sum_{n\geq 2}n^{2}\,|\varepsilon_{n}|^{2},\qquad\gamma_{q}=\alpha\,(q-1),$ (17)
where $\alpha=\mathcal{O}(1)$ is a non-universal constant determined by the microscopic embedding. Combining Eqs. (16) and (17), the total quadratic action for the soft reparametrization modes becomes
$S^{(2)}_{\rm tot}=2\pi\bar{C}T\sum_{n\geq 2}\Big{[}n(n^{2}-1)+\gamma_{q}n^{2}\Big{]}\,|\varepsilon_{n}|^{2},$ (18)
where the first term $n(n^{2}-1)$ is the standard Schwarzian spectrum after fixing the $SL(2,\mathbb{R})$ zero modes, while the second term $\gamma_{q}n^{2}$ represents the leading $q$-deformation effect. Physically, the additive $\gamma_{q}n^{2}$ term modifies the dispersion of the soft reparametrization modes. While the Schwarzian spectrum $n(n^{2}-1)$ grows cubically with $n$, the $q$-deformation introduces a quadratic contribution that enhances the stiffness of low-$n$ modes, most prominently the $n=2$ component. As a result, the $q$-deformation shifts the balance between universal Schwarzian fluctuations and non-Gaussian corrections inherited from the microscopic dynamics. The Matsubara propagator for the reparametrization (soft) modes is defined as the frequency-space two-point function
$D(i\omega_{n})\equiv\langle|\varepsilon_{n}|^{2}\rangle=\frac{\displaystyle\int\mathcal{D}\varepsilon\;|\varepsilon_{n}|^{2}\,e^{-S^{(2)}_{\rm tot}[\varepsilon]}}{\displaystyle\int\mathcal{D}\varepsilon\;e^{-S^{(2)}_{\rm tot}[\varepsilon]}}.$ (19)
Here $\varepsilon_{-n}=\varepsilon_{n}^{*}$ and the functional measure factorizes into independent integrals over the complex Fourier modes $\varepsilon_{n}$ for each $n\geq 2$. Because the total action (18) is quadratic and diagonal in Fourier space, the functional integral factorizes into a product of complex Gaussian integrals. For a single complex variable $z\in\mathbb{C}$ the standard identities
[…]
imply
$\frac{\int d^{2}z\,|z|^{2}e^{-A|z|^{2}}}{\int d^{2}z\,e^{-A|z|^{2}}}=\frac{1}{A}.$
Applying this to each Fourier mode with quadratic kernel $2\pi\bar{C}TA_{n}$ (where $A_{n}\equiv n(n^{2}-1)+\gamma_{q}n^{2}$) yields the exact Matsubara propagator in the Gaussian approximation
$D(i\omega_{n})=\langle|\varepsilon_{n}|^{2}\rangle=\frac{1}{2\pi\bar{C}T}\,\frac{1}{\,A_{n}\,}=\frac{1}{2\pi\bar{C}T}\;\frac{1}{n(n^{2}-1)+\gamma_{q}n^{2}}.$ (20)
Analytic continuation $i\omega_{n}\to\omega+i0^{+}$ and interpolation to the continuum $n\sim\omega/(2\pi T)$ yield the retarded propagator $D^{R}(\omega)=D(i\omega_{n}\to\omega+i0^{+})$, whose imaginary part defines the spectral density
[…]
with a non-universal prefactor $\kappa_{b}=\mathcal{O}(1)$. The deformation introduces an effective scale
$T_{0}\equiv\frac{\alpha}{2\pi}(q-1)\Lambda\;\;\propto\;\;\gamma_{q}T,$ (22)
where $\Lambda\sim 2\pi T$ is the thermal UV cutoff. The emergent Tsallis scale $T_{0}$ is quantitatively anchored by the Raman low-frequency slope, whose ratios across doping levels yield consistent values around $100\text{K}$, as shown in the Appendix. Equation (21) shows that the $q$-deformation regularizes the infrared behavior of the soft modes by shifting the effective bath temperature to $T+T_{0}$. In the undeformed limit $q\to 1$, one recovers the canonical Schwarzian bath with Ohmic spectral density $A_{\varepsilon}(\omega)\propto|\omega|/T$ [39].
Having established the quadratic dynamics of the soft reparametrization modes and their Matsubara propagator, we now turn to their coupling to microscopic matter fields. In particular, under a boundary reparametrization $\tau\!\to\!f(\tau)$, a fermion transforms as
$\psi(\tau)\;\longrightarrow\;[f^{\prime}(\tau)]^{1/2}\,\psi\!\big{(}f(\tau)\big{)}.$ (23)
Expanding $f(\tau)=\tau+\varepsilon(\tau)$ to linear order in the soft mode $\varepsilon(\tau)$ and gauge-fixing the $SL(2,\mathbb{R})$ zero modes, one finds a derivative coupling between the soft mode and the local fermionic energy density $T_{\tau\tau}=\psi^{\dagger}i\partial_{\tau}\psi$:
$S_{\rm int}=g\int_{0}^{\beta}d\tau\;\varepsilon^{\prime}(\tau)\,\psi^{\dagger}(\tau)\,i\partial_{\tau}\psi(\tau),$ (24)
where $g=\mathcal{O}(1)$ in natural units ($\hbar=k_{B}=1$). Fourier transforming to frequency space, $\varepsilon(\tau)=\sum_{\omega}\varepsilon_{\omega}e^{-i\omega\tau}$ and $\psi(\tau)=\sum_{\varepsilon}\psi_{\varepsilon}e^{-i\varepsilon\tau}$, the vertex for soft-mode exchange takes the form
$\Gamma(\varepsilon,\omega;\varepsilon-\omega)=g\,\varepsilon+\mathcal{O}(\omega),$ (25)
where $\varepsilon$ denotes the fermionic Matsubara frequency (measured from the chemical potential) and $\omega$ the soft-mode frequency. The leading dependence $\propto\varepsilon$ reflects that reparametrizations couple directly to the stress tensor $T_{\tau\tau}$, while the subleading terms $\mathcal{O}(\omega)$ represent higher derivative corrections suppressed at low frequencies. The linear-in-frequency coupling structure in Eq. (25) naturally leads to an Ohmic-type spectral density at low energies, consistent with the universal dissipative behavior expected from soft reparametrization modes [40, 41, 42, 43].
When a fermion couples to a bosonic bath such as the Schwarzian soft mode, the fermionic self-energy in real time can be computed by summing all contour contractions according to Wick’s theorem. At second order, the retarded self-energy is given by the standard Keldysh expression
$\Sigma^{R}(\varepsilon)=i\int\!\frac{d\omega}{2\pi}\,\Big{[}\Gamma^{2}D^{K}(\omega)G^{R}_{0}(\varepsilon-\omega)+\Gamma^{2}D^{R}(\omega)G^{K}_{0}(\varepsilon-\omega)\Big{]},$ (26)
where $G^{R/K}_{0}$ are the bare fermionic Green’s functions and $D^{R/K}$ are the retarded and Keldysh soft-mode propagators. Note that the bare fermionic Green’s functions in the time domain are $G_{0}^{R}(t)=-i\,\theta(t)\,\langle\{c(t),c^{\dagger}(0)\}\rangle$ and $G_{0}^{A}(t)=i\,\theta(-t)\,\langle\{c(t),c^{\dagger}(0)\}\rangle$. Using the fluctuation–dissipation relations
$D^{K}(\omega)=\big{[}D^{R}(\omega)-D^{A}(\omega)\big{]}\coth\!\left(\frac{\omega}{2T}\right),\qquad G^{K}_{0}(\varepsilon)=\big{[}G^{R}_{0}(\varepsilon)-G^{A}_{0}(\varepsilon)\big{]}\tanh\!\left(\frac{\varepsilon}{2T}\right),$ (27)
[…]
$\Gamma(\varepsilon)=\int_{0}^{\infty}\!\frac{d\omega}{2\pi}\;\underbrace{\big{[}g^{2}\,A_{\varepsilon}(\omega)\big{]}}_{K(\omega)}\big{[}n_{B}(\omega)+n_{FD}(\omega-\varepsilon)\big{]},$ (28)
[…]
$K(\omega)=\frac{\kappa}{\bar{C}}\,\frac{|\omega|}{T_{0}+T},\qquad\kappa=\frac{g^{2}\kappa_{b}}{2\pi},$ (29)
which makes explicit that the effective interaction strength scales inversely with the renormalized Schwarzian stiffness $\bar{C}$.
After integration, we obtain the compact form
$\Gamma(T)=\frac{\pi^{2}\tilde{g}^{2}}{4}\,\frac{T^{2}}{T_{0}+T},$ (30)
which precisely matches the Golden-Rule scaling widely used in phenomenological analyses of strange-metal transport. Corrections arising from finite-$\delta$ deformations or higher-order interactions contribute only subleading terms, such as $\mathcal{O}(T/T_{0})$ or logarithmic corrections at strong coupling. The controlled Keldysh derivation clarifies that the Ohmic kernel $K(\omega)\propto\omega/(T_{0}+T)$ originates from the soft-mode spectral density and derivative coupling in Eq. (24). The temperature dependence $T^{2}/(T_{0}+T)$ in Eq. (30) follows from integrating over thermal phase space up to $\omega\sim T$ and is consistent with both perturbative and phenomenological treatments of electron–Schwarzian scattering.
Substituting the soft-mode scattering rate into the semiclassical relaxation-time formula,
$\sigma_{xx}\simeq e^{2}\int d\epsilon\,(-\partial_{\epsilon}f)\,N(\epsilon)\,v_{x}^{2}(\epsilon)\,\tau_{\rm tr}(\epsilon),\qquad\tau_{\rm tr}(\epsilon)\sim\Gamma^{-1}(\epsilon),$ (31)
one obtains a temperature-dependent conductivity. Using the result of Eq. (30), the corresponding resistivity takes the form
$\rho_{xx}(T)\equiv\sigma_{xx}^{-1}(T)\;\propto\;\frac{T^{2}}{T_{0}+T}.$ (32)
This scaling law reveals a smooth crossover: at low temperatures ($T\ll T_{0}$), the resistivity follows the Fermi-liquid form $\rho\!\sim\!T^{2}$, while at high temperatures ($T\gg T_{0}$) it becomes linear, $\rho\!\sim\!T$, as commonly observed in strange metals. The interpolation $\rho(T)\!\sim\!T^{2}/(T_{0}+T)$ therefore provides a natural explanation for the empirical crossover from quadratic to linear resistivity, arising microscopically from Schwarzian soft-mode scattering [28, 33].More broadly, experimental studies have suggested that the linear-in-$T$ resistivity slope may correlate with superconducting scales in certain cuprates and pnictides [44, 45, 46, 47], although such connections lie beyond the scope of the present analysis.
A caveat in the above transport estimate is that the Schwarzian soft modes are intrinsically 0+1 dimensional (they possess no spatial momentum), whereas electrical conduction is a spatial transport phenomenon. To make contact with conductivity one therefore needs an additional ingredient that converts local (time-only) scattering into momentum/charge relaxation. In this work we adopt the physically transparent assumption that the Schwarzian modes act as a local (momentum-independent) dissipative bath for spatially extended fermions; the resulting single-particle inelastic rate $\Gamma(\varepsilon)$ computed above is then supplemented by a separate momentum-relaxing channel (e.g. Umklapp processes) that ensures finite DC conductivity. Under the further assumption that vertex corrections are subleading (or that the scattering is approximately angle-independent on the Fermi surface), the transport time can be taken as $\tau_{\rm tr}(\varepsilon)\!\sim\!\Gamma^{-1}(\varepsilon)$, and the semiclassical Drude estimate applies, yielding $\rho(T)\propto T^{2}/(T_{0}+T)$.
Hall angle from the memory matrix- The memory-matrix formalism provides a controlled framework for transport in systems with slow or nearly conserved operators, where standard Boltzmann theory fails [48, 49, 50]. The central idea is that dc conductivities are determined by the relaxation of such slow operators through their couplings to a dissipative bath. For strange metals, this bath is supplied by Schwarzian soft modes, whose Ohmic spectrum ensures smooth low-frequency behavior and introduces an additional channel for energy dissipation. Concretely, for a basis of slow operators $\{A_{\alpha}\}$ the conductivity tensor reads
$\sigma_{ij}=\sum_{\alpha\beta}\chi_{J_{i}A_{\alpha}}\,\big{[}M(0)+N\big{]}^{-1}_{\alpha\beta}\,\chi_{A_{\beta}J_{j}},$ (33)
where $M$ encode the dissipative couplings to the bath, $\chi_{\alpha\beta}=\langle A_{\alpha};A_{\beta}\rangle$ the static susceptibilities. This expression applies equally to Fermi liquids and non–Fermi liquids. In the Fermi-liquid case, $M$ reduces to the standard quasiparticle collision integral, recovering the textbook Kubo–bubble results. $N$ is the antisymmetric block induced by the Lorentz force, with $N_{P_{x}P_{y}}=-N_{P_{y}P_{x}}=\omega_{c}\chi_{PP}$ and $\omega_{c}=eB/m^{\ast}$. For completeness, we rederive the dc conductivity and the quadratic magnetoresistance of a conventional Fermi liquid in Appendices A and B. The key quantities are the relaxation rates
$\Gamma_{Q}\equiv\frac{M_{QQ}(0)}{\chi_{QQ}},\qquad\Gamma_{P}\equiv\frac{M_{PP}(0)}{\chi_{PP}},$
which characterize dissipation in the energy-like channel $Q$ and in the momentum channel $P$, respectively. When $Q$ is chosen as the current-like soft mode, Eq. (33) reduces to
$\sigma_{xx}=\frac{\chi_{JJ}}{\Gamma_{Q}},\qquad\rho_{xx}=\frac{\Gamma_{Q}}{\chi_{JJ}}.$ (34)
[…]
$M_{QQ}(0)\propto T\qquad\Rightarrow\qquad\rho_{xx}\propto T.$ (35)
Physically, the soft reparametrization modes relax energy efficiently without relaxing total momentum, giving rise to a Planckian-like linear resistivity channel. For the momentum block,
[…]
In a continuum model one has $\dot{P}_{x}=0$, so momentum relaxation arises solely from translation-breaking terms such as static disorder $V(\mathbf{r})$ and lattice umklapp. For a weak periodic potential $V(\mathbf{r})=\sum_{\mathbf{G}}U_{\mathbf{G}}e^{i\mathbf{G}\cdot\mathbf{r}}$, the coefficients $U_{\mathbf{G}}$ are Fourier components of the lattice potential at reciprocal vectors $\mathbf{G}$. The operator $\rho_{\mathbf{G}}=\sum_{\mathbf{k},\sigma}c^{\dagger}_{\mathbf{k}+\mathbf{G},\sigma}c_{\mathbf{k},\sigma}$ transfers electronic density by momentum $\mathbf{G}$, while $G_{x}$ denotes the $x$-component of $\mathbf{G}$, quantifying the amount of momentum lost along the transport direction. The leading memory-matrix element controlling momentum relaxation then reads
[…]
where $\chi^{R}$ is the retarded density–density correlator of the interacting system. In clean correlated metals the dominant temperature dependence arises from electron–electron umklapp scattering at finite $\mathbf{G}$. Phase-space integrals over two quasiparticles near the Fermi surface yield the familiar Fermi-liquid factor
$\int d\epsilon_{1}d\epsilon_{2}\,\epsilon_{1}\epsilon_{2}\,(-\partial_{\epsilon_{1}}f)(-\partial_{\epsilon_{2}}f)\;\propto\;T^{2},$ (38)
so that
$M_{PP}(0)\propto T^{2},\qquad\Gamma_{P}=\frac{M_{PP}}{\chi_{PP}}\propto T^{2}.$ (39)
Hence momentum relaxation is parametrically slower than energy relaxation, consistent with conventional large-angle umklapp scattering. Because $Q$ couples strongly to the current but weakly to the Lorentz-force channel, while $P$ couples directly to the Lorentz block $N$, the transport coefficients separate naturally:
$\rho_{xx}\propto\Gamma_{Q}\sim T,\qquad\cot\theta_{H}\simeq\frac{M_{PP}}{\omega_{c}\chi_{PP}}\propto T^{2}.$ (40)
This dichotomy encapsulates the two-lifetime phenomenology of strange metals. Deviations from $\cot\theta_{H}\propto T^{2}$ in electron-doped cuprates [51] and ruthenates [52] reflect orbital selectivity and multi-band scattering, but the separation of channels remains a robust organizing principle [53].
The canonical scaling of the Hall angle, $\cot\theta_{H}\propto T^{2}$, observed in optimally doped cuprates and several strange metals, has been widely interpreted as evidence for distinct transport and Hall scattering rates [53]. This behavior, however, is not universal. In electron-doped cuprates, such as $\mathrm{Pr_{2-x}Ce_{x}CuO_{4-\delta}}$ [51], and in layered ruthenates including $\mathrm{Sr_{2}RuO_{4}}$ and $\mathrm{Sr_{3}Ru_{2}O_{7}}$ [52], clear departures from the $T^{2}$ law have been reported. The coexistence of multiple Fermi-surface sheets with distinct orbital characters leads to orbital-dependent scattering lifetimes that evolve differently with temperature and magnetic field. In $\mathrm{Sr_{2}RuO_{4}}$, for instance, the $\alpha$, $\beta$, and $\gamma$ bands derived from different Ru $4d$ orbitals exhibit markedly different coherence scales and gap-like features [52]. Such orbital-selective effects effectively renormalize the temperature dependence of the transport tensors, producing an apparent breakdown of the simple scaling. These observations point to the necessity of introducing a material-dependent cutoff scale $T_{0}$, which encodes band-structure and orbital complexity within a unified transport framework.
Within the memory-matrix framework, once Schwarzian soft modes are incorporated, the separation between energy relaxation ($\Gamma_{Q}\sim T$) and momentum relaxation ($\Gamma_{P}\sim T^{2}$) arises naturally. This accounts for the canonical TT–linear resistivity and $T^{2}$ Hall angle in single-band strange metals, while also providing a platform to understand deviations in multiband systems. Orbital selectivity modifies the effective relaxation rates but does not alter the fundamental two-channel structure that distinguishes $\rho_{xx}$ and $\cot\theta_{H}$.
Having established how distinct scattering channels control the longitudinal and transverse responses, we now turn to the role of an external magnetic field. The same formalism offers a natural route to magnetotransport, and in particular to the microscopic origin of the scale-invariant, linear-in-$B$ magnetoresistance observed in quantum-critical strange metals. The scattering rate is given by
$\Gamma(\varepsilon)=\int_{0}^{\infty}\frac{d\omega}{2\pi}\;K(\omega;B)\big{[}n_{B}(\omega)+n_{FD}(\omega-\varepsilon)\big{]},$ (41)
[…]
$A(\omega;B)\simeq\frac{\kappa_{b}}{\bar{C}}\frac{|B|\,|\omega|}{T_{0}+T},$ (42)
then the kernel $K(\omega;B)=g^{2}(B)A(\omega;B)$ directly yields
$\Gamma(0)=\frac{\kappa\pi}{8\bar{C}}\frac{|B|\,T^{2}}{T_{0}+T},$ (43)
so that at fixed $T$ the longitudinal resistivity scales linearly with field, $\rho_{xx}\propto|B|$. This mechanism captures key experimental observations in quantum-critical strange metals. In BaFe2(As1-xPx)2, near the QCP the scattering rate follows [54]
$\Gamma\approx\sqrt{(\alpha k_{B}T)^{2}+(\eta\mu_{B}B)^{2}},\quad\frac{\eta}{\alpha}\approx 1,$ (44)
demonstrating nearly identical thermal and magnetic energy scales. Similarly, in overdoped La2-xSrxCuO4, scale-invariant linear-in-$B$ resistivity is observed up to $80~\mathrm{T}$ [55], with slope $\beta\sim\alpha k_{B}/\mu_{B}$, consistent with Eq. (43). Together, these results are consistent with a Planckian-limited scaling,
$\Gamma(T,B)\propto\max\{\alpha k_{B}T,\eta\mu_{B}B\},$ (45)
operating across a fan-shaped quantum-critical region. This scaling suggests that field-enhanced electron–soft mode scattering provides a natural microscopic route to linear magnetoresistance, in line with experimental observations in pnictides [54], cuprates [55], and the broader phenomenology of Planckian dissipation [12, 13]. The memory-matrix analysis unifies the two-lifetime phenomenology ($\rho_{xx}\sim T$, $\cot\theta_{H}\sim T^{2}$) with the scale-invariant magnetoresistance ($\rho_{xx}\sim|B|$), underscoring the central role of soft-mode physics in non-Fermi-liquid transport.
Discussion and Conclusion-Starting from the quantum Tsallis distribution, we have developed a unified framework for strange-metal transport. The $q$-deformation naturally introduces soft reparametrization modes whose dynamics can be captured by a Schwarzian-type action. These collective modes provide an intrinsic channel for energy relaxation, interpolating smoothly between Fermi-liquid and non-Fermi-liquid regimes.
Within this framework, the resistivity evolves from the familiar quadratic temperature dependence at low temperatures to a linear behavior at high temperatures, reproducing the hallmark of strange metals. The memory-matrix analysis further clarifies that longitudinal transport is controlled by energy relaxation into soft modes, while the Hall response is dictated by the slower momentum relaxation associated with weak translation-symmetry breaking. This separation leads to two distinct lifetimes, a feature widely observed experimentally. The framework also accounts for the crossover from linear-in-field magnetoresistivity to the quadratic regime, in agreement with Kohler scaling analyses.
Beyond transport, the present approach suggests new directions for understanding the interplay between strange-metal behavior and high-temperature superconductivity. Soft-mode fluctuations associated with the Tsallis deformation may enhance pairing interactions or reshape the electronic density of states, thereby influencing the superconducting transition. Future work should focus on quantitative calculations of the gap, critical temperature, and coherence length, and on systematic comparisons with experimental observables such as resistivity slopes, Hall angles, and magnetoresistance across different families of high-$T_{c}$ materials.
Acknowledgments- We thank Xiao Hu, Sang-Jin Sin, Yan Liu, Matteo Baggioli and Kui Jin for helpful discussions. We would also like to express our gratitude to colleagues in the "Eight Immortals Crossing the Sea" WeChat group for their engaging discussions on topics beyond physics. This work was partially supported by NSFC, China (Grant Nos. 12275166 and 12311540141).
Appendix A DC Conductivity of Fermi liquid
We present the standard Kubo-bubble derivation and show its reduction to the semi-classical (relaxation-time) form used in the main text.
A.1 Kubo bubble and analytic continuation
The optical conductivity $\sigma_{xx}(\omega)$ of an interacting electron system can be obtained from the retarded current–current correlation function [50]:
$\Pi^{R}_{xx}(\omega)=-i\int_{0}^{\infty}dt\,e^{i\omega t}\langle[J_{x}(t),J_{x}(0)]\rangle,$ (46)
where $J_{x}=-e\sum_{\mathbf{k}}v_{x}(\mathbf{k})c^{\dagger}_{\mathbf{k}}c_{\mathbf{k}}$ is the $x$-direction current operator and $v_{x}(\mathbf{k})=\partial\varepsilon_{\mathbf{k}}/\partial k_{x}$ is the group velocity for band dispersion $\varepsilon_{\mathbf{k}}$. Within the bubble (one-loop) approximation in the Matsubara formalism, the correlation reads
$\Pi_{xx}(i\Omega_{n})=-e^{2}T\sum_{i\omega_{m}}\sum_{\mathbf{k}}v_{x}^{2}(\mathbf{k})G(\mathbf{k},i\omega_{m})G(\mathbf{k},i\omega_{m}+i\Omega_{n}),$ (47)
where $G(\mathbf{k},i\omega_{m})$ is the full single-particle Green’s function, and $T$ is the temperature (natural units $\hbar=k_{B}=1$ are used). The spectral representation,
$G(\mathbf{k},i\omega_{m})=\int_{-\infty}^{\infty}\frac{d\epsilon}{2\pi}\frac{A(\mathbf{k},\epsilon)}{i\omega_{m}-\epsilon},$ (48)
introduces the spectral function $A(\mathbf{k},\omega)=-2\,\mathrm{Im}\,G^{R}(\mathbf{k},\omega)$ and the Fermi distribution
$f(\epsilon)=\frac{1}{e^{(\epsilon-\mu)/T}+1}.$ (49)
After analytic continuation $i\Omega_{n}\to\omega+i0^{+}$, the retarded correlator becomes
$\Pi^{R}_{xx}(\omega)=-e^{2}\sum_{\mathbf{k}}v_{x}^{2}(\mathbf{k})\int_{-\infty}^{\infty}\frac{d\epsilon}{2\pi}\frac{f(\epsilon)-f(\epsilon+\omega)}{\omega+i0^{+}}A(\mathbf{k},\epsilon)A(\mathbf{k},\epsilon+\omega).$ (50)
The optical conductivity is then
$\sigma_{xx}(\omega)=\frac{1}{\omega}\,\mathrm{Im}\,\Pi^{R}_{xx}(\omega),$ (51)
where the spectral function $A(\mathbf{k},\omega)$ encodes interaction effects and quasiparticle lifetimes.
A.2 DC limit and the quasi-particle approximation
Taking the dc limit $\omega\to 0$ in Eq. (50), the difference in Fermi distributions can be expanded as
$f(\epsilon)-f(\epsilon+\omega)\simeq-\omega\,\partial_{\epsilon}f(\epsilon),$
yielding
$\sigma_{xx}=e^{2}\sum_{\mathbf{k}}v_{x}^{2}(\mathbf{k})\int_{-\infty}^{\infty}\frac{d\epsilon}{2\pi}\big{(}-\partial_{\epsilon}f(\epsilon)\big{)}A^{2}(\mathbf{k},\epsilon),$ (52)
where $\partial_{\epsilon}f(\epsilon)$ restricts contributions to energies within a thermal window $\sim T$ around the Fermi level. Here, $A(\mathbf{k},\epsilon)$ is the single-particle spectral function, and $v_{x}(\mathbf{k})=\partial\varepsilon_{\mathbf{k}}/\partial k_{x}$ is the group velocity along the $x$-direction.
Near the Fermi surface, the spectral function is well approximated by a Lorentzian form
$A(\mathbf{k},\epsilon)=\frac{1}{\pi}\frac{\Gamma(\mathbf{k},\epsilon)}{(\epsilon-\xi_{\mathbf{k}})^{2}+\Gamma^{2}(\mathbf{k},\epsilon)},$
where $\xi_{\mathbf{k}}=\varepsilon_{\mathbf{k}}-\mu$ is the band energy measured from the chemical potential $\mu$, and
[…]
is the quasiparticle scattering rate determined by the imaginary part of the retarded self-energy $\Sigma^{R}$. In the sharp-quasiparticle limit ($\Gamma\ll T$), the squared spectral function simplifies to
$A^{2}(\mathbf{k},\epsilon)\simeq\frac{2\pi}{\Gamma(\mathbf{k},\epsilon)}\,\delta(\epsilon-\xi_{\mathbf{k}})+\mathcal{O}(1).$
Substituting this expression into Eq. (52) and integrating over energy and momentum yields the relaxation-time formula
$\boxed{\sigma_{xx}\simeq e^{2}\int d\epsilon\;(-\partial_{\epsilon}f(\epsilon))\,N(\epsilon)v_{x}^{2}(\epsilon)\tau_{\rm tr}(\epsilon)},$ (53)
where
$\tau_{\rm tr}(\epsilon)\simeq\frac{1}{2\Gamma_{\rm tr}(\epsilon)}$ (54)
is the transport relaxation time and
$N(\epsilon)=\sum_{\mathbf{k}}\delta(\epsilon-\xi_{\mathbf{k}})$
is the single-particle density of states (DOS). Equation (53) is equivalent to the Boltzmann relaxation-time approximation when vertex corrections are negligible, capturing the essential semiclassical physics of momentum relaxation [50].
In a Landau Fermi liquid, the low-energy scattering rate scales as
$\Gamma_{\rm tr}(\epsilon,T)\propto(\epsilon-\mu)^{2}+(\pi T)^{2},$ (55)
reflecting phase-space restrictions for electron–electron scattering. Inserting this into Eq. (53), and evaluating near the Fermi level ($\epsilon\simeq\mu$), one obtains
$\rho_{xx}(T)\equiv\frac{1}{\sigma_{xx}(T)}\propto T^{2},$ (56)
the hallmark quadratic temperature dependence of resistivity in a conventional Fermi liquid [2]. Deviations from this $T^{2}$ law signal non-Fermi-liquid behavior, often arising near quantum critical points or in strongly correlated regimes.
Appendix B Fermi-liquid limit: Quadratic magnetoresistance
In the memory-matrix framework, the formal expression
$\sigma_{ij}=\sum_{\alpha\beta}\chi_{J_{i}A_{\alpha}}\big{[}M(0)+N\big{]}^{-1}_{\alpha\beta}\chi_{A_{\beta}J_{j}}$
remains valid for both Fermi-liquid (FL) and non-Fermi-liquid (non-FL) scenarios. The distinction lies in the microscopic $B$-dependence of the memory-matrix entries.
For a conventional Fermi liquid, taking the minimal momentum basis $\{P_{x},P_{y}\}$ and assuming field-independent susceptibilities, the dominant momentum-relaxation element admits a weak-field expansion
$M_{PP}(B)=M_{0}+\alpha B^{2}+O(B^{4}),$
with no linear-in-$|B|$ contribution. Inverting the $2\times 2$ block then yields the standard conductivity formulas, leading to the quadratic magnetoresistance
$\rho_{xx}(B)=\rho_{0}+\Delta\rho(B),\qquad\Delta\rho(B)\propto(\omega_{c}\tau_{\rm tr})^{2}\propto B^{2},$
which arises when the magnetic field does not generate new low-frequency scattering channels. This includes impurity- or phonon-dominated elastic scattering (field-independent $\tau_{\rm tr}$) and conventional FL inelastic scattering with $\tau_{\rm tr}^{-1}\propto T^{2}$. The expansion is valid in the weak-field limit $\omega_{c}\tau_{\rm tr}\ll 1$ and continuous spectral densities. Departures from this behavior—such as Landau quantization or field-enhanced soft modes—lead to non-FL magnetoresistance, including the linear-in-$|B|$ response discussed in the main text.
More generally, within the same framework the momentum-relaxation element may be written as
$M_{PP}(B)=M_{0}+c|B|+\alpha B^{2}+\cdots,$ (57)
where the coefficient $c$ encodes scattering enhanced by soft modes. The longitudinal resistivity then takes the form
$\rho_{xx}(B)=\frac{M_{PP}(B)}{\chi_{JP}^{2}}+\frac{\omega_{c}^{2}\chi_{PP}^{2}}{\chi_{JP}^{2}M_{PP}(B)},\qquad\omega_{c}=\frac{eB}{m^{\ast}},$ (58)
which, upon expansion for small to moderate fields, yields
$\rho_{xx}(B)\simeq\rho_{0}+\frac{c}{\chi_{JP}^{2}}|B|+\left(\frac{e^{2}\chi_{PP}^{2}}{m^{\ast 2}\chi_{JP}^{2}M_{0}}+\frac{\alpha}{\chi_{JP}^{2}}\right)B^{2}+O(B^{3}),\qquad\rho_{0}=\frac{M_{0}}{\chi_{JP}^{2}}.$ (59)
The crossover field $B^{\ast}\sim cM_{0}/\beta$, with $\beta=e^{2}\chi_{PP}^{2}/(m^{\ast 2}\chi_{JP}^{2})$, separates the linear-in-$|B|$ regime ($|B|\ll B^{\ast}$) from the quadratic FL regime ($|B|\gg B^{\ast}$). Experimentally, Kohler plots collapse $\Delta\rho/\rho_{0}$ versus $(\omega_{c}\tau_{\rm tr})^{2}$ in the FL regime, while a robust linear $|B|$ slope signals non-FL scattering channels [57, 49, 56, 54, 55]. This demonstrates that the memory-matrix approach interpolates continuously between quadratic FL magnetoresistance and linear non-FL magnetoresistance within a single microscopic framework.
Appendix C Experimental estimation of $T_{0}$
The aim of this section is to determine the emergent Tsallis scale $T_{0}$ that governs non-Fermi liquid scattering in correlated materials. This scale enters the parametrization $q(T)=1+T/T_{0}$ of the Tsallis distribution and sets the strength of soft-mode fluctuations responsible for marginal transport. Experimentally, $T_{0}$ can be estimated from Raman scattering, where the low-frequency slope of the electronic continuum directly reflects the quasiparticle scattering rate. Defining the slope as $S\equiv\left.\tfrac{dI(\Omega)}{d\Omega}\right|_{\Omega\to 0^{+}}$, one can compare two samples at the same temperature and construct the ratio
$R_{2}(T)=\frac{S_{\rm sample}(T)}{S_{\rm ref}(T)}.$ (60)
Within the Tsallis parametrization $q(T)=1+T/T_{0}$, this ratio admits the inversion formula
$T_{0}(T)=\frac{T}{R_{2}(T)-1},$ (61)
so that experimental Raman slopes immediately yield an estimate of $T_{0}$.
For cuprates, Chang et al. [1] demonstrated that at the critical doping $p_{c}\approx 0.19$, the electronic self-energy evolves from an incoherent regime ($\mathrm{Im}\,\Sigma\sim\text{const.}$) to a coherent Fermi-liquid regime ($\mathrm{Im}\,\Sigma\sim T^{2}$). The Raman $B_{1g}$ low-frequency slope tracks this crossover: it is $\sim 0.5$–1 eV-1 near optimal doping ($p\approx 0.16$) and increases significantly for $p>p_{c}$ in the strange-metal regime. Complementary thermodynamic measurements by Talluri et al. [2] show that the electronic specific heat coefficient $\gamma(T)=C_{\rm el}/T$ is suppressed for $p<p_{c}$, but converges to a linear-in-$T$ form above $p_{c}$, reinforcing the identification of $p_{c}$ as the strange-metal–pseudogap boundary. In iron-based superconductors (Ba(Fe1-xCox)2As2, BFCA), Raman $B_{1g}$ slopes reported by Gallais and Paul [3] and Jost et al. [4] display a similar enhancement near the nematic quantum critical point. Using Eq. (61), slope ratios such as $(n_{\rm dop}=0.04;0.02)$ and $(n_{\rm dop}=0.24;0.07)$ give $T_{0}\simeq 166$–233 K, while for Bi2212 cuprates, comparing $(n_{\rm dop}=0.21;0.19)$ yields $T_{0}\simeq 154$–200 K, consistent across independent datasets.
Taken together, these results show that in both iron-based and cuprate superconductors the extracted Tsallis scale $T_{0}$ clusters around $\mathcal{O}(10^{2})\,\text{K}$. This convergence demonstrates that soft-mode fluctuations of comparable strength underpin marginal scattering, Raman continua, specific heat anomalies, and ultimately high-$T_{c}$ superconductivity across distinct material families. The unified scale $T_{0}$ suggests a universal energy scale for quantum critical fluctuations driving high $T_{c}$ superconductivity, linking Raman, specific heat, and transport anomalies across cuprates and iron-based superconductors.